### v1

# MindShift: Leveraging Large Language Models for Mental-States-Based Problematic Smartphone Use Intervention

Abstract.
Problematic smartphone use negatively affects physical and mental health. Despite the wide range of prior research, existing persuasive techniques are not flexible enough to provide dynamic persuasion content based on users’ physical contexts and mental states. We first conduct a Wizard-of-Oz study (N=12) and an interview study (N=10) to summarize the mental states behind problematic smartphone use: boredom, stress, and inertia. This informs our design of four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. We leverage large language models (LLMs) to enable the automatic and dynamic generation of effective persuasion content. We develop MindShift, a novel LLM-powered problematic smartphone use intervention technique. MindShift takes users’ in-the-moment physical contexts, mental states, app usage behaviors, users’ goals & habits as input, and generates high-quality and flexible persuasive content with appropriate persuasion strategies. We conduct a 5-week field experiment (N=25) to compare MindShift with baseline techniques. The results show that MindShift significantly improves intervention acceptance rates by 17.8-22.5% and reduces smartphone use frequency by 12.1-14.4%. Moreover, users have a significant drop in smartphone addiction scale scores and a rise in self-efficacy. Our study sheds light on the potential of leveraging LLMs for context-aware persuasion in other behavior change domains.
Problematic smartphone use, persuasion, large language model, mental model
ccs: Human-centered computing Empirical studies in HCI
1. Introduction
In recent years, the ubiquitous presence of smartphones has increased people’s reliance on digital devices. This trend frequently leads to problematic smartphone usage behaviors (Hwang et al., 2012), especially among adolescents and young adults (Lee et al., 2014). Prior studies suggest that problematic smartphone usage can detrimentally affect people in various areas such as efficiency (leading to diminished academic or work performance (Arefin et al., 2018; Duke and Montag, 2017)), physical well-being (resulting in decreased sleep and activity levels (Clayton et al., 2015; Lapointe et al., 2013; Xu et al., 2019, 2021)), and mental health (manifesting as anxiety and depression (Hartanto and Yang, 2016; Harwood et al., 2014; Xu et al., 2022a)). Many individuals have recognized their problematic smartphone usage and sought to reduce their over-reliance on smartphones (Hiniker et al., 2016a; Ko et al., 2015).
There has been a plethora of academic research and commercial products that provide just-in-time (JIT) smartphone use interventions that intervene at the moment when problematic use is taking place (Nahum-Shani et al., 2018). Existing intervention techniques can be summarized into four categories based on their degree of enforcement: (1) Feedback: offering insights on phone usage patterns, enhancing user awareness about their smartphone habits (Whittaker et al., 2016; Kim et al., 2016; Collins et al., 2014; Andone et al., 2016); (2) Reminders: aiming to counter immediate phone indulgence and promote self-reflection through pop-up notifications (Hiniker et al., 2016a; Park and Gweon, 2015); (3) Interaction friction: raising the effort needed to use the phone, thereby reducing its allure by introducing tasks like typing (Park et al., 2018; Okeke et al., 2018; Xu et al., 2022b); (4) Lockout: disabling the user’s phone access for a specified duration(Kim et al., 2019a; Ko et al., 2015; Löchtefeld et al., 2013; Kim et al., 2019b).
Figure 1. Overview of MindShift. When users exhibit problematic phone usage, MindShift actively collects data on phone usage behavior, physical contexts, and mental states, and uses the customized persuasion strategies we designed along with users’ goals and habits, to generate prompts. Then, the large language model will generate a persuasion message. Finally, the persuasion will show up in users’ phones to encourage more mindful usage.
However, we notice a few gaps among existing intervention techniques. One major issue is the specific moment of intervention. Most existing intervention techniques are time-based, such as simply upon app launch (Xu et al., 2022b; Kim et al., 2019b), or after hitting the upper limit of usage duration (Hiniker et al., 2016a; Kim et al., 2019a; Park et al., 2018). However, these methods ignore an important factor: users’ smartphone use purpose. Users’ smartphone use purposes can be generally divided into instrumental use (i.e., for specific goals such as working or intentional relaxation) and habitual use (i.e., ritualistic behavior, without a clear goal, such as passive social media content consumption) (Rubin, 2009; Hiniker et al., 2016b). Recent research suggests intervention should mainly focus on habitual use since it often gives people a lower sense of meaningfulness (Lukoff et al., 2018). Thus, an intervention technique should consider not only the use time but also the smartphone usage purpose.
Furthermore, we identify a further gap in the intervention content. Existing JIT intervention techniques rarely take users’ in-the-moment contexts (e.g., current environment, mental states) into account to achieve context-aware intelligent intervention (Thomas Craig et al., 2021). This is particularly important for smartphone persuasion intervention, one of the reminder-based techniques that show messages designed to convince users to stop using smartphones. The persuasion content is the key to the effectiveness of intervention. Although some intervention techniques are designed to be context-aware (e.g., an intervention is introduced when users are around a specific location (Kim et al., 2017; Ko et al., 2015)), they are simply rule-based and cannot adapt to the dynamic changing physical environment. In addition, no prior work considers users’ mental state, which is another important aspect of user contexts.
To address these gaps, we first conducted a Wizard-of-Oz study (N=12), followed by an interview study (N=10) to better understand users’ mental states during problematic phone usage. Focusing on habitual usage, we summarized three major mental states to address: boredom, stress, and inertia. Leveraging the Dual Systems Theory (Hofmann et al., 2009) and the ERG (Existence, Relatedness, and Growth) Theory (Caulton, 2012), we proposed a set of four persuasion strategies and built the mapping between these strategies to the three mental states. These strategies include 1) understanding, 2) comforting, 3) evoking, and 4) scaffolding habits.
Based on our persuasion strategies, we designed and implemented MindShift (Figure 1), a new JIT intervention technique that can provide dynamic, personalized persuasion content based on user contexts. MindShift leverages Large Language Models’ (LLMs) strong capability in commonsense comprehension and natural language generation (Chowdhery et al., 2022; Min et al., 2021). Specifically, we employ LLMs to generate proper and effective persuasive content based on phone usage behavior, physical contexts, mental states, goals and habits with appropriate persuasion strategies we design.
To evaluate the effectiveness of MindShift, we conducted a 5-week field experiment to deploy our intervention to 25 participants. We compared MindShift against the basic notification-based intervention that asked users to reflect and report their smartphone usage purpose. Moreover, to better evaluate the effect of the mental states factor, we also compared MindShift against a simplified version that excludes the mental states factor from the LLM-based content generation.
Our study results indicate that MindShift significantly outperforms the baseline method by 17.8-22.5% on the intervention acceptance rate. MindShift also reduces overall app opening frequency by 12.1-14.4% and app usage duration by 2.4-9.8%. Specifically, MindShift leads to a significant reduction in habitual proportions by 6.8%. Comparing the two versions of MindShift, including the mental states factor in the persuasion content generation can significantly enhance the acceptance rate by 8.1%. The subjective report data show that participants have a significant decrease (25.8-34.7%) in the smartphone addiction scale (SAS) score and a significant increase (10.4-10.7%) in the self-efficacy scale score after using MindShift. Participants also expressed a strong willingness to continue using this effective intervention approach.
The contribution of our paper can be summarized as follows:
  1. (1)
    We conducted a Wizard-of-Oz study and an interview study to investigate the mental states that end-users experience during smartphone habitual use. Our study highlights three major mental states (boredom, stress, and inertia), leading us to design four persuasion strategies grounded in the Dual Systems Theory and the ERG Theory: Understanding, Comforting, Evoking, and Scaffolding Habits.
  2. (2)
    We created MindShift, a novel persuasive intervention technique that leveraged an LLM to generate dynamic and personalized persuasion content based on appropriate persuasion strategies, users’ phone usage behavior, physical contexts, mental states, goals and habits.
  3. (3)
    We conducted a field experiment by deploying MindShift. Our results indicate that our new technique could significantly improve intervention acceptance rates and reduce smartphone use. Users’ subjective feedback also corroborates these observations, demonstrating the effectiveness of MindShift.
2. Related work
In this section, we first introduce the definition of smartphone problematic use and habitual phone use (Sec. 2.1). We then briefly overview existing intervention techniques (Sec. 2.2). Close to our work, we introduce the behavior change persuasion techniques and their relationship with the boost of LLMs (Sec. 2.3). Finally, we cover the theoretical foundations and mental models of problematic smartphone use (Sec. 2.4).
2.1. Smartphone Problematic Use
Many studies have explored the definition of problematic smartphone use, which can be broadly classified into two categories. The first category defines whether users exhibit addictive behaviors toward their phones. Some studies assessed addictive behavior by measuring the level of user dependence on smartphones through questionnaires, such as the Smartphone Addiction Scale (SAS) and the Smartphone Addiction Inventory (SPAI) (Lin et al., 2014; Kwon et al., 2013). The second category defines whether a specific instance of phone use is problematic. Recent research suggests that problematic smartphone use is determined not only by excessive use but also by the purpose and content of use in specific situations (Harwood et al., 2014; Prasad et al., 2021). Studies have pointed out that users’ purposes for phone use can be categorized into (1) habitual use that is performed unconsciously and ritualistically usually without a specific goal (Pinder et al., 2018), and (2) instrumental use with a specific task or goal in mind (Lukoff et al., 2018). Existing research suggests that habitual use should be the primary target for intervention (Morris et al., 2018; Lukoff et al., 2018; Pinder et al., 2018). In this paper, we combine these two definitions. We use SAS to measure the level of addictive smartphone usage behavior. We also distinguish users’ phone use intention and focus on intervening habitual use.
2.2. Problematic Smartphone Use Intervention
Existing problematic smartphone use intervention techniques can be roughly divided into four categories based on their level of enforcement: The first category provides users with information about their behavior (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a; Collins et al., 2014; Apple, 2021; Lyngs et al., 2022; Nwagu, 2023), requiring users to view it themselves to increase awareness of phone usage. The second actively sends notifications and reminders to users to provoke their reflection such as reminding users of their daily goals (Hiniker et al., 2016a), sending them the results of previous decision-balancing exercises (Park and Gweon, 2015), or informing their usage time (Apple, 2021). This category presents text to users and, therefore, serves as a persuasion. The third involves increasing the difficulty of using the phone and intentionally slowing down user interactions to suppress the desire to use it. This can include requiring users to enter random numbers or engage in self-reflection (Park et al., 2018; Xu et al., 2022b), keeping the phone vibrating continuously (Okeke et al., 2018). The fourth is particularly forceful by directly locking the users’ apps or phone for a specific duration (APP, 2021; Kim et al., 2019a; Ko et al., 2015).
There are concerns about these methods’ ability to strike an optimal balance between user experience and effectiveness. The first two intervention categories (i.e., feedback-based self-awareness and notification reminders) rely on the user’s self-control and can be easily ignored, leading to limited effectiveness (Whittaker et al., 2016; Kim et al., 2016; Hiniker et al., 2016a). While the other two intervention types (i.e., increasing usage difficulty and locking out) are relatively more effective, research has shown that users often struggle to maintain compliance due to a diminished user experience (Kim et al., 2019a). Once the restrictions are removed, users often relapse and revert to problematic use behavior in a short period of time (Kim et al., 2019a; Xu et al., 2022b). To bridge the gap, we hope to develop a method that ensures a good user experience while boosting intervention effectiveness.
2.3. Persuasion for User Behavior Change and Large Language Models
Persuasion is a psychological approach designed to influence attitudes, beliefs, or behaviors(Cialdini and Cialdini, 2007). Language is the most commonly used means of persuasion, and this approach leverages facts, emotional appeals, and so on in a message to achieve its goal. Its effectiveness has been shown in multiple fields, such as advertising to encourage consumers to buy a product(Bernritter et al., 2017), supporting mental health such as coping with stress (Morris et al., 2018; Orji and Moffatt, 2018), and managing physical health such as reducing snacking behavior (Kaptein et al., 2012). For smartphone use intervention, persuasion usually appears as reminders, such as leveraging the user’s usage time in a template format (Apple, 2021; Inc, 2023), or some thought-provoking statements (Xu et al., 2022b). Prior work has suggested that the personalization of persuasive content can enhance the effectiveness of intervention (Choe et al., 2017; Harrison et al., 2015; Kaptein et al., 2012). Some studies pointed out that the effectiveness of static, unchanging intervention methods would decrease over time, while dynamic intervention methods with diverse intervention timings or contents (even appearing randomly) can improve effectiveness (Kovacs et al., 2018).
The recent emergence of Large Language Models (LLMs), such as ChatGPT(OpenAI, 2022) and PaLM(Chowdhery et al., 2022), has made vast progress in the generation of personalized and diverse content. Recent studies have explored various health applications supported by LLMs, such as health information seeking (Yunxiang et al., 2023), mental health support (Lamichhane, 2023; Kumar et al., 2023), personal health coaching (Wei et al., 2023; Montagna et al., 2023), and health education (Montagna et al., 2023; Kung et al., 2023). Recent research in the HCI community has also explored LLMs’ potential in public health interventions (Jo et al., 2023). However, there is a gap in leveraging LLMs for a real-time, context-aware behavior intervention. In this work, we aim to leverage LLMs’ ability to dynamically generate personalized persuasion content to help users reduce their problematic phone usage.
2.4. Theoretical Background and Mental Models of Smartphone Use
Understanding what leads to problematic phone use is essential for the design of effective persuasion strategies. The Dual Systems Theory (Hofmann et al., 2009) has been used to explain the phone usage patterns (Pinder et al., 2018). This theory divides human cognitive activities into two types: System 1 (fast, intuitive, unconscious) and System 2 (slow, analytical, conscious). Problematic smartphone use is typically driven by System 1 (Hagger, 2016), as it mainly involves unconscious, rapid responses and is easy to be guided by instant gratification. Recent research suggests that two key factors contribute to the failure of users to act on their goals: (i) limited ability of System 2 control; and (ii) fluctuations caused by emotional states and fatigue (Lyngs et al., 2019). These findings suggest that mental states play a significant role in habitual smartphone usage(Pinder et al., 2018). While previous research has identified external contexts that may trigger the habitual use of smartphones (Tran et al., 2019), limited research has explored users’ mental states as a factor. In our work, we pinpoint the major mental states linked to habitual use and propose corresponding persuasion strategies for these mental states.
3. Mental States of Habitual Smartphone Use and Persuasion Strategies
To better understand the mental states of users’ smartphone use and guide the design of our intervention system, we first conducted a Wizard-of-Oz (WoZ) study, followed by a semi-structured interview study (Sec. 3.1). We summarize the results in Sec. 3.2. Based on our initial findings, we devise four persuasion strategies (Sec. 3.3).
3.1. Exploratory Wizard-of-Oz & Semi-structured Interview Studies
To identify particular smartphone usage behaviors requiring targeted interventions and further develop persuasive intervention content material, we first conducted a WoZ study in the wild, involving 12 end-users (6 females and 6 males, aged 18-28), followed by a semi-structured interview with another 10 users (5 females and 5 males, aged 18-29).
3.1.1. Wizard-of-Oz Study
We developed a chatbot system for smartphones that tracks user app activity. First, we asked participants to select apps for intervention, adding them to a blacklist. The chatbot then would send persuasive messages to the participants upon opening a blacklisted app.
At the beginning of the study, participants were told that it was an automatic chatbot instead of a human to reduce the observation effect (Kazdin, 1982). In reality, when participants opened a blacklisted app, a human experimenter would receive an email notification. Based on participants’ smartphone usage data, the experimenter designed and delivered persuasive messages aiming to reduce smartphone usage via the chatbot. Each evening, researchers carried out a brief daily interview to probe (1) why participants use smartphones, (2) factors affecting their acceptance of interventions, and (3) evaluation of the effectiveness of the persuasive messages. Our WoZ study lasted five days for each participant. At the study’s conclusion, we informed participants that the chatbot was operated by a human experimenter at the back end.
3.1.2. Semi-structured Interview Study
We also carried out a semi-structured interview study with a different participant group. During the semi-structured interview, we asked participants to recollect instances of problematic smartphone use. Next, we asked them to discuss their mental states during those instances. We followed the participants’ lead during the interview. All interviews were audio recorded and transcribed. Three researchers independently examined the transcriptions and then collectively coded them through multiple iterations until they reached a consensus.
3.2. Main Takeaways about Problematic Smartphone Use
First, interventions for problematic smartphone use should target habitual usage. Our WoZ study delineates two primary types of smartphone usage: instrumental and habitual, which corroborates the classification of prior research  (Lukoff et al., 2018). We find that only habitual smartphone use warrants intervention. Providing interventions during instrumental use often leads to user dissatisfaction. Notably, relaxation emerges as a crucial form of instrumental use, where participants deliberately use their phones to unwind or reward themselves after intense work or study. Intervening at such times is considered intrusive and inappropriate. This classification aligns with the Dual Systems Theory. In smartphone interactions, instrumental use relies on conscious decision-making (System 2), while habitual use is more instinctive (System 1). Hence, interventions should primarily target habitual use, a conclusion also supported by earlier studies  (Lukoff et al., 2018).
Second, the effectiveness of interventions depends on users’ mental states, personal goals, and contextual information. This is consistent with the literature, suggesting that a shift away from System 2 is due to emotional fluctuations and the absence of defined goals and intentions (Lyngs et al., 2019). We experimented with different persuasive message content during our WoZ study. We find that when we incorporate users’ mental states factor, which is inferred by their physical contexts and app usage patterns, into persuasive message content, participants have an increased acceptance rate of the intervention. In addition, we find emphasizing users’ personal goals can enhance the interventions’ effectiveness. For instance, sending persuasive messages to participants such as, “When you find yourself with idle time, consider engaging in meaningful activities such as reading, writing, or drawing” proves effective when users are in an idle state and aimed for self-improvement. Our findings are supported by prior studies linking habitual smartphone use to specific mental states  (Anderson and Wood, 2021; Bayer and LaRose, 2018).
Third, semi-structured interviews reveal three primary mental states connected to habitual smartphone use: boredom, stress, and inertia.
  • Boredom is an affective state characterized by low arousal and dissatisfaction due to insufficient stimulation  (Fisherl, 1993; Mikulas and Vodanovich, 1993). The WoZ and interview studies identify common scenarios leading to boredom: (1) when the task at hand is too simple, lacking a balance between skill and challenge, (2) lack of interest in the current activity, and (3) devoid of any engaging activities during idle moments, which can trigger cognitive biases like the fear of missing out and promote habitual use  (Bozaci, 2020; Smith, 2015; Wolniewicz et al., 2020).
  • Stress refers to cognitive and behavioral reactions to unpredictable and unmanageable stimuli (Koolhaas et al., 2011). Participants frequently use smartphones because they experience (1) heightened anxiety when work demands exceed their abilities and (2) uncertainty about whether something would have a positive outcome. This aligns with increasing evidence links smartphone use to perceived stress (Chiu, 2014; Samaha and Hawi, 2016; Vahedi and Saiphoo, 2018).
  • Inertia, in our context, refers to a psychological resistance that makes users reluctant to change their current activity state. It is similar to the idea illustrated by literature such as emotional inertia (Kuppens et al., 2010) or decision-making inertia (Alós-Ferrer et al., 2016). Our research shows that participants commonly use their phones habitually to avoid changing into a new activity state from an idle state (e.g., checking the phone before starting to work). Unlike stress or boredom, inertia does not elicit overt negative emotions but impedes the shift to the next task.
Moreover, we find that users’ mental states are either triggered by specific activities or have no specific reason. Therefore, we categorize scenarios based on whether users are engaged in specific activities (“engaging in activities”, “not engaging in activities”). The first category (“engaging in activities”) denotes mental states triggered by ongoing or upcoming activities. For instance, activities that lack challenge and engagement, or activities that involve repetitive actions might induce boredom. In contrast, the second category (“not engaging in activities”) has less clear causes, often linked to broader life situations. For example, a lack of enthusiasm for life may cause boredom, while challenges adapting to setbacks and uncertainties can cause stress. Hence, combining the three mental state factors with whether or not engaging in activities factor, we categorize habitual phone use into six distinct scenarios (see the gray part of Figure 2).
Figure 2. Summary of Users’ Mental States and Corresponding Persuasion Strategies Based on WoZ and Interview Studies. In the 3 x 2 table (three mental states: boredom vs. stress vs. inertia, and two activities: engage in vs. not engage in), each color denotes the persuasion strategy applicable in this scenario.
3.3. Persuasion Strategies Design
Based on the Dual Systems Theory and the mental states we summarized in Sec 3.2, we propose four distinct persuasion strategies: Understanding, Comforting, Evoking, and Scaffolding Habits (see the colorful part in Figure 2).
Our strategy design is inspired by the theory of Existence, Relatedness, and Growth (ERG). Developing from Maslow’s Hierarchy of Needs, ERG theory further summarizes human motivation into three levels: (i) the physiological and safety basic needs for existence, (ii) the social needs for feeling related and accepted, and (iii) the need to grow and self-actualize. The theory has been applied to workplaces to increase productivity and job satisfaction (Arnolds and Boshoff, 2002; Caulton, 2012).
Maintaining System 2 relies on cultivating motivation for meaningful activities. Thus, we map strategies to the relatedness level and growth level of the ERG theory in Figure 3. At the level of relatedness, Understanding and Comforting aim to empathize with users’ emotions, offering support and empowering users to manage System 2. At the growth level, Evoking reminds users of their personal development goals, and Scaffolding Habits guides them in replacing habitual smartphone use with activities conducive to self-fulfillment, thereby turning awareness into action.
Figure 3. Mapping of Persuasion Strategies to the ERG Theory (Existence, Relatedness, and Growth) and the Dual Systems Theory. According to the Dual Systems Theory, there is a competitive relationship between System 1 (habitual smartphone use) and System 2 (meaningful activity), much like being placed on a scale. To make System 2 heavier than System 1, weights are added—Relatedness needs and Growth needs (the second and third levels of ERG Theory). To support users’ Relatedness needs, we design two persuasion strategies: Understanding and Comforting. To motivate Growth needs, we design two other persuasion strategies: Evoking and Scaffolding Habits.
3.3.1. Understanding
Understanding is a critical strategy to motivate users at the Relatedness level. Past literature suggests that seeking understanding is a coping mechanism (Compas et al., 2001), and chatbots’ empathetic expressions are favored over emotionally neutral advice (Liu and Sundar, 2018). Therefore, Understanding is one of our strategies and covers all mental states in Figure 2. For instance, this example shows how we integrate understanding into the persuasive content intervention: ”Hi, I know that sometimes you may feel bored and lacking interest. It’s okay, this is a very normal feeling. Everyone goes through such times.”
3.3.2. Comforting
Comforting aims to comfort users who are experiencing emotional fluctuations, which cause them to shift away from System 2. Prior studies suggest that coping with boredom should focus on meaningfulness  (Nett et al., 2010). For example, we design a persuasive message to say ”Hey, I know some things might seem a bit boring, but sometimes we need to find the fun in them. Have you ever thought that completing this task would bring you closer to your goals?” In addition, encouragement, humor, acceptance, and wishful thinking (Compas et al., 2001) can be used to cope with stress by lightening the uncontrollable and unpredictable nature of stressors. For example, ”Don’t worry, you have the capability to complete the task! Believe in yourself, and the outcome will pleasantly surprise you.” In the state of inertia without explicit negative emotions, Comforting is not employed.
3.3.3. Evoking
We find evoking specific work-related goals is a compelling, persuasive technique based on the WoZ study. Literature suggests that goals and values held in working memory are important for people to sustain System 2 (Lyngs et al., 2019) and are closely related to growth motivation. Thus, Evoking considers users’ goals (e.g., maintaining physical health, achieving academic success) for designing persuasive strategies. For example, ”Hi! I know you want to play with your phone, but completing tasks is crucial for your IELTS! Keep going, and you’re one step closer to a high score!” This strategy is applied only to scenarios that which users are engaged in activities, because they may lack work-related goals when are not engaged in activities.
3.3.4. Scaffolding Habits
Last, we encourage users to develop alternative beneficial habits to excessive smartphone use, aligning with their personal value and growth need (Lyngs et al., 2019; Pinder et al., 2018). By pre-identifying users’ preferred habits and considering variables like location and time of habitual smartphone use, we suggest appropriate substitutes to assist users in Scaffolding Habits. For example, ”Hi, why not use this moment to memorize vocabulary instead of using your phone? It can help you learn a language and achieve your goals faster!”
Through our Wizard of Oz (WoZ) and interview studies, we target habitual use for intervention and pinpoint crucial mental states that drive habitual usage. We then apply the Dual Systems Theory and ERG Theory to develop four targeted persuasion strategies. This informs the design of our MindShift system.
4. MindShift DESIGN
Building on top of the persuasion strategies we propose in the previous section, we introduce the design of our intervention system, including persuasive content generation (what content to intervene, Sec.  4.1), the interaction flow (how to intervene, Sec.  4.2), and the intervention moment (when to intervene, Sec.  4.3).
4.1. What Content to Intervene: LLM-Powered Persuasive Content Generation
Figure 4. Prompt Templates Used to Persuasive Generate Content for Smartphone Usage Intervention with GPT-3.5. The word slots (in the top box) represent different categories of information to be filled in based on the user’s current situation. The color indicates the mapping between the slots and the input prompt. The small words under each slot explain the source of the content. The input prompt consists of four parts: (1) Task Setup; (2) Description of the Current Contexts; (3) Prompt Optimization (to improve language quality and reduce harmful content); and (4) Description of Persuasion Strategy (as introduced in Section 3 and Figure 2). The LLM output shows the persuasion example of four strategies when the user’s mental state is “stressful, engaging in an activity”.
One of the key aspects that influence the effectiveness of our intervention is the persuasion content design. We propose a set of important content factors and couple them with an LLM for high-quality and intelligent content generation. We test both GPT-3.5 and GPT-4 and choose to adopt GPT-3.5 as our target LLM to strike the balance between the content generation quality and the speed.  [^1] GPT-4 introduces a long lag and negatively impacts user experience. Moreover, as we introduce below, the prompts we use as the input for GPT-3.5 include general information that is not individually identifiable. However, we do acknowledge the privacy risk of our method. We will have more discussion in Section 8 Our prompt design incorporates four factors, arranged in a sequence from 1 to 4, as illustrated by purple text in Figure 4. The combined prompt subsequently serves as input to LLMs to generate persuasive content for tailored intervention.
(1) Task Setup: As shown in the left top box in the ”Input Prompt to LLM” in Figure 4, Task Setup includes ¡Background¿ module, which provides GPT-3.5 with the global instructions.
(2) Description of the Current Contexts: To make each generated content contextually relevant and personalized, it is necessary to describe the user’s current contexts (see the left bottom box in the Input prompt in Figure 4). This part includes three modules: the user’s physical context (¡User Data¿ module), mental state (¡User Mental State¿ module), and their goals (¡User Goals¿ module). (a) The ¡User Data¿ module describes the user’s real-time physical context, including the current time, location, habitual phone usage duration, and the time elapsed since the last habitual phone check. (b) The ¡User Mental State¿ module includes users’ input from their devices regarding activities and negative emotions. The prompt for this module is selected from the mental state definition (the gray part of Figure 2). (c) Finally, the ¡User Goals¿ module describes what the user values and plays a crucial role in the generation of Evoking strategies. Specific use goals are collected during the initialization (Step 1 in Figure 5). The elements mentioned above are represented as word slots (enclosed in brackets in the input prompt), where the actual users’ context information can be inserted.
(3) Prompt Optimization: To improve the quality and effectiveness of the content generated by GPT, we carefully crafted the prompt according to OpenAI’s official guidelines (OpenAI, 2023). We acknowledge sometimes LLMs can generate harmful, offensive, or biased texts (Zhuo et al., 2023; Hartvigsen et al., 2022). To address the ethical concerns, we employed an iterative prompt design process to ensure that the persuasive content is appropriate. Initially, one researcher conducted a preliminary design and invoked the GPT-3.5 API to generate 30 sets of content using artificial data. Subsequently, two other researchers rated the satisfaction level of the generated content (scoring from 1 to 5). This process is iterated multiple rounds until the average satisfaction score is above 4. This process strives to create content that is both appropriate and engaging, aligning with our persuasive strategy. We note that this process cannot fully address the ethical concerns. We discuss the ethical concerns and LLM’s use risk in Section 8. The middle top box in Figure 4 shows our final optimization prompt. Prompt Optimization includes two modules: (a) The ¡Notes¿ module clarifies restrictions; (b) The ¡Output Format¿ module guarantees the correct format of the output.
(4) Description of Persuasion Strategy: To ensure each generated content aligned with our proposed strategies, we provide a short description of each strategy, as indicated by the middle bottom box in Figure 4. This description is based on our design in Section 3 (also see Figure 2). Based on users’ different in-the-moment mental states, we will select the corresponding strategies. Note that for the scaffolding habits strategy, we also input a habit selected from users’ initialization. The way we obtain the mental states and specific habits is mentioned in Section 4.2.
Figure 5. Interaction Flow. Users first set their value and app list (Step 1). When opening a black-listed app, users need to first self-report their phone use intent (Step 2). If the intent is habitual use, the app asks them to report their current mental state (Step 3). After that, a corresponding persuasion shows (Step 4).
4.2. How to Intervene: Interaction Flow
We design a four-step user interaction flow for MindShift, as depicted in Figure 5.
Step 1: Initialization. When initiating MindShift, we ask users to set their values in four categories: (Step 1 in Figure 5) career, health, life, and hobbies. For each category, users are asked to fill in both their goals and actions. This process serves two purposes: first, it provides the necessary user-specific information for our Evoking strategy; second, it enables the creation of personalized habits in the Scaffolding Habits persuasion strategy. After that, users are asked to set a blacklist of apps. Launching apps from this list will trigger intervention.
Step 2: Intention report. Based on our findings in Section 3, interventions are only necessary when users are habitually using their phones. To collect users’ intentions, we employ a self-reporting approach (Step 2 in Figure 5). When an intervention is triggered, users are presented with three options to choose from: habitual use, instrumental use, and relaxation. Only when users select ’habitual use’, the intervention will proceed. Since reporting intent itself is an intervention, driving users to reflect on their behavior, we add an exit button at the bottom of the interface to offer an exit from the app. It is noteworthy that detecting habitual use is beyond the scope of this paper. We envision future work on contextual sensing and user intent detection that can automate this process. We have more discussion in Section 8.
Step 3: Current mental state report. The mental state is a key factor that triggers users’ habitual phone use. Unlike physical context, which can be automatically detected, current mental state detection techniques are not mature and robust enough. Therefore, we ask users to self-report (Step 3 in Figure 5). Based on the mental states listed in Section 3.2, we propose two single-choice questions for users to report their mental state: (1) whether they are engaged in activities (”Yes” or ”No”), and (2) whether they currently have any negative feelings, including options of ”Stress”, ”Boredom”, ”None” (i.e., inertia), or ”Other Negative Feelings”. The ”Other Negative Feelings” option is used to capture other cases that offer choices that do not cover. Users can input in a text box next to the radio button.
Step 4: Persuasion. Based on our intervention content design in Figure 4 and our persuasion strategy, we leverage the power of GPT-3.5 to generate persuasive messages. The messages are displayed in a pop-up window. Below the message, users can choose to either quit the app or continue using it (the bottom of Step 4 in Figure 5). Additionally, users can also provide optional feedback by giving a thumbs up or down to the intervention content.
Furthermore, for the Scaffolding Habits strategy, it is essential to link between users’ own habits and their specific use contexts. Therefore, we implement a user participation mechanism. The interface for scaffolding habits intervention has an additional habit item along with an edit button (the upper interface of Step 4 in Figure 5). The habit item represents a system-generated suggestion based on the user’s current context and initial settings. Users have the control to edit and define their desired habits. Once submitted, the modified habit will be recommended the next time when users are in the same context.
4.3. When to Intervene: Intervention Trigger Mechanism
As in Figure 2, each mental state allows for multiple persuasion strategies. Therefore, we devise a simple procedure. After determining the mental state and narrowing down the specific strategies, we first pick one strategy and generate an intervention message. Then, we loop over other different strategies and show new strategy messages every two minutes until the users leave the app. After looping over all appropriate strategies under this mental state, the intervention will stop. Users only need to report (i.e., Step 2 and 3 in Figure 5) the first time they open a specific app after each screen unlock. This design is based on two considerations: 1. We assume that the user’s initial opening usually reflects their subsequent usage intentions. 2. Requiring users to report multiple times when switching between apps can lead to a poor user experience.
5. System Implementation
We built an Android application to instantiate our design. Our system consists of a client and a server.
5.1. Client-Side Implementation
The client-side is an Android app that implements all the features of our design. We use accessibility services to detect the opening and closing of apps on the phone. The client is also responsible for collecting and uploading data to the server, including the screen’s off and unlock status, the name, opening and closing times of the application, and location data obtained through the Amap API (Amap, 2023). To prevent the accidental killing of the accessibility service and ensure compliance, we also add a background service in the app to check the status of the accessibility service every 5 minutes. If it detects that the accessibility service is terminated, the client will notify the server to send an email to a researcher. The researcher can remind the user to turn on the accessibility service again to ensure the integrity of the data.
5.2. Server-Side Implementation
The server side is responsible for generating persuasive content tailored to the user’s context through four key tasks. By performing these tasks, the server side ensures that the intervention is personalized, contextually relevant, and delivered in a timely manner.
(1) User Data Computation: The server first computes the user data that needs to be used in the word slots from the app data uploaded by the client, including the total habitual usage time of the phone and the time of the last habitual opening.
(2) Habit Selection: Next, the server selects a habit mostly matched with the current user’s mental state, location (i.e., the specific building), and time (i.e., the hour of the day) from users’ initialization (i.e., Step 1 in Figure 5). To create a strong association between habit and context, as long as users do not thumb down or modify it, the same habit will be recommended in the same context. More details can be seen in Step 4 of 4.2.
(3) Strategy Counterbalance: To ensure that the number of times each strategy appears in each mental state is roughly the same, the server then counterbalances the strategy order in the prompt mentioned in Section 4.2.
(4) Content Generation: After obtaining the user contexts, habits, and persuasion strategy, the server uses the OpenAI GPT-3.5 API to generate appropriate intervention content. We adopted a streaming, character-by-character generation approach, allowing the persuasive content to start being displayed within 2 seconds.
6. Field Experiment
To evaluate the effectiveness of MindShift, we conducted a 5-week field experiment. In this section, we introduce experimental design, participant recruitment, and experiment procedure.
6.1. Baseline Intervention Methods
As MindShift is one of the first persuasion intervention systems leveraging an LLM to generate dynamic persuasion content, there are no comparable systems other than the traditional persuasion techniques. We compare MindShift against a persuasive reminder baseline (also as a pop-up window), one of the most commonly adopted intervention methods in commercial apps (Inc, 2023; Apple, 2021). To ensure the fairness of the comparison, the persuasion baseline is designed to be the same as the intent report step, as illustrated in Step 2 in Figure 5. We name this intervention as Baseline.
Moreover, in order to evaluate the effectiveness of the mental states and persuasion strategies proposed in Section 3, we further design a simplified version of MindShift by removing the mental states and persuasion strategies from the prompt design. The content generation will solely be based on users’ physical contexts and their app usage behavior. We name this group as MindShift-Simple. In total, we have three intervention methods to compare: Baseline, MindShift-Simple, and MindShift. Table 1 shows the comparison of these three intervention techniques.
Table 1. Comparison of Three Intervention Techniques
Intervention Methods Characteristics
Intent Report LLM-powered Persuasion Mental-State-based Persuasion Strategies
Baseline $\checkmark$
MindShift-Simple $\checkmark$ $\checkmark$
MindShift $\checkmark$ $\checkmark$ $\checkmark$
6.2. Experiment Design
We adopt a within-subject design, with intervention techniques as independent variables (Baseline, MindShift-Simple, and MindShift). We design a 5-week field experiment. To measure users’ everyday phone usage behavior, the first week is set as the Baseline stage, followed by two weeks of one MindShift version and another two weeks of the other version. We counter-balance the order of MindShift-Simple and MindShift.
Our evaluation metrics include various aspects: (1) Intervention acceptance rate. We measure the percentage of times users accept the intervention and quit the blacklist app use when interventions are shown. (2) Intervention thumb-up rate. For MindShift-Simple, and MindShift that show persuasion content, users can provide feedback by pressing buttons (thumb-up or thumb-down, see Figure5 step 4 interface). (3) App usage behavior, which includes both app opening frequency and usage duration. (4) Subjective reports. At the beginning of the study and at the end of each intervention session, we distributed the Smartphone Addiction Scale (SAS) (Kwon et al., 2013) and the self-efficacy Scale (Schwarzer and Jerusalem, 1995). In addition, at the end of the study, we also conducted a brief semi-structured interview to gather user experiences and feedback on our intervention techniques. These metrics cover both the objective and subjective measures of the interventions.
6.3. Participants
We sent out recruitment material on social media platforms. We included a screening survey aiming to identify potential participants who showed signs of smartphone addiction and the willing to reduce their smartphone use. Specifically, besides basic demographics, we included four questions selected from the SAS and self-efficacy questionnaires. We also included questions about the willingness to reduce smartphone use, the extent of habitual phone use, future plans for the next five weeks, and a screenshot of phone usage time of the last week. We excluded users who (1) did not show signs of smartphone addiction (SAS sub-score ¡ 15), or (2) were unwilling to reduce smartphone use, or (3) had a total weekly usage time of less than 20 hours, or (4) those who had special plans such as long-term travel in the next five weeks (which may shift their phone usage patterns).
We received a total of 42 responses. We recruited 31 participants after the screening process. 6 participants voluntarily dropped out during the study. For the remaining participants, we divided them into groups according to counterbalanced intervention orders. We conducted a Kruskal-Wallis test analysis to ensure that groups had no significant difference in the SAS scores and self-efficacy scores. In the end, 25 participants completed the entire study (females=13, males=12, age=22±2 years). Among them, 17 participants were undergraduates, 5 were graduate students, and 3 were professionals.
6.4. Experiment Procedure
After all the participants signed the consent form, we held a 20-minute onboarding session online to familiarize participants with the research process and introduce the Android application. We explained in detail the meanings of each selection in the report interfaces (Step 2 and 3 in Figure 5). After the meeting, participants filled out the first SAS and self-efficacy questionnaires. The app was then deployed for a 5-week field experiment.
Before users started using one of two versions of MindShift (MindShift and MindShift-Simple), we provided participants with a tutorial explaining the three mental states (i.e., boredom, stress, and inertia) that they need to report. This ensured that they had an accurate and consistent understanding of the mental states, which in turn improved the accuracy and reliability of the reported data. At the end of the experiment, participants received a compensation of \$50 for their time.
7. RESULTS
During the five-week study, we collect 50,815 minutes of restricted app usage duration, and 54,467 restricted app opening events (7539, 23,769, 21,994, and 835 for habitual use, instrumental use, relax, and quit). We analyze the quantitative data collected via app and the qualitative data collected via questionnaires and interviews.
7.1. Intervention Acceptance Rate
The effectiveness of a persuasion strategy is directly measured by the rate of successful prevention of user engagement with the targeted application (i.e., intervention acceptance rate). We compare the overall acceptance rate (Sec. 7.1.1), acceptance rate for generated persuasion content (Sec. 7.1.2), and thumb-up rate (Sec. 7.1.3). Moreover, we also conduct a detailed analysis of the acceptance rate across strategies (Sec. 7.1.4), mental states and activities (Sec. 7.1.5).
7.1.1. Overall Acceptance Rate
Figure 6. Overall Acceptance Rate
MindShift and MindShift-Simple increase the overall acceptance rate significantly and MindShift achieve best. We first investigate the overall acceptance rate during the three intervention sessions, the total number of quit times divided by the total number of app visits (excluding instrumental uses and relaxation). Figure 6 shows the comparison, MindShift (35.2±24.1%) has a higher acceptance than MindShift-Simple (30.5±23.6%) and Baseline (12.7±17.4%). These differences are statistically significant according to a Friedman test ($\chi^{2}$(2) = 13.69, p ¡ .01). Three post hoc Wilcoxon signed-rank tests, corrected with Holm’s sequential Bonferroni procedure, indicate that MindShift vs. Baseline (V = 52, p ¡ .01) and MindShift-Simple vs. Baseline (V = 38, p ¡ .001) are statistically significantly different, but that MindShift vs. MindShift-Simple is not significant (V = 129, n.s.).
7.1.2. Acceptance Rate for Generated Persuasion Content
(a) Overall Persuasion Acceptance Rate
(b) Persuasion Acceptance Rate Grouped by Round
Figure 7. Persuasion Acceptance Rate
MindShift has a significantly higher persuasion acceptance rate than MindShift-Simple. The overall acceptance rate includes two parts: exiting when reporting intention and exiting after seeing the persuasion content. To narrow down the comparison between MindShift and MindShift-Simple, we focus on the intervention acceptance rate during the persuasion stage, as the two intervention techniques’ differences lie only in the persuasive content. As shown in Figure 6(a), MindShift achieves higher persuasion acceptance (20.1±20.2%) than MindShift-Simple (12.0±15.0%) and a paired-samples t-test shows that MindShift was statistically significantly higher (t = -2.21, p¡0.05).
Moreover, as we introduce in Section 4.3, every persuasion intervention could consist of 1 to 4 rounds, depending on which mental state participants are in and which stage participants leave the app. Therefore, we further compare the persuasion acceptance rates between MindShift and MindShift-Simple across different rounds. The results show that MindShift outperforms MindShift-Simple at each round ($\Delta_{\textit{Round1}}$=6%, $\Delta_{\textit{Round2}}$=14.7%, $\Delta_{\textit{Round3}}$=10.2%, $\Delta_{\textit{Round4}}$=11%) as indicated in Figure 6(b). Moreover, we observe a decreasing trend of the acceptance rate of MindShift-Simple, while MindShift does not show such a trend. We conduct a paired-samples t-test between the two intervention techniques in each round. The results indicate that, except for the first round, rounds 2 (p¡0.05), 3 (p¡0.001), and 4 (p¡0.001) all exhibit that MindShift achieves significantly higher acceptance rate.
7.1.3. Thumb-up Rate of Interventions
(a) Feedback Proportion
(b) Acceptance Rate Grouped by Persuasion Strategies
Figure 8. Feedback and Strategy Acceptance Rate
MindShift has a significantly higher thumb-up rate. Users can give feedback in the persuasion interface. As depicted in Figure 7(a), 6.8% of interventions in MindShift receives thumb-up while MindShift-Simple receives only 2.2%. A paired-samples t-test shows that significant differences (p¡0.05) are observed for the thumb-up rate, but no significant differences (p=0.22) for the thumb-down rate between the two techniques. This indicates that MindShift aligns better with users’ preferences.
7.1.4. Acceptance Rate across Different Strategies
[…]
(a) Strategy Acceptance Rate Grouped by Mental State. Significant differences compared to no strategy (green bar) are indicated above the X coordinate axis, while differences among strategies are below.
(b) Strategy Acceptance Rate Grouped by Activity. The same annotation method as grouped by mental state.
Figure 9. Strategy Acceptance Rate across Different Mental States and Activities
7.1.5. Strategy Acceptance Rate across Different Mental States and Activities
As we show in Figure 2, each mental state has a different strategy mapping. Therefore, we also sought to derive insights regarding which strategies are most effective for users under different mental states (Figure 8(a)) and activities (Figure 8(b)). Friedman test and post hoc Wilcoxon signed-rank tests are employed to investigate the influence of different strategies on acceptance rate.
[…]
For activity levels: When not engaging in activity, Comforting is significantly more effective than MindShift-Simple ($\Delta_{\textit{}}$=5.6%, p ¡ 0.05). When engaging in activity, there are no significant differences in all the strategies compared to MindShift-Simple.
7.2. App Usage Behavior
We then investigate the influence of the intervention on participants’ app usage behavior. Overall, participants have less app usage frequency and duration when using MindShift and MindShift-Simple, especially in habitual usage.
7.2.1. Overall Usage Behavior
We count the number of app opening attempts for restricted apps. Figure 9(a) presents the opening frequency under three intervention techniques. MindShift-Simple (63.6±9.2) and MindShift (65.3±9.5) have lower opening frequency than Baseline (74.3±9.7). Compared to Baseline, MindShift reduces by 12.1% usage duration while MindShift-Simple reduces by 14.4%. However, a Friedman test does not show significance.
We also measure restricted app usage duration which is another important factor for phone overuse. As can be seen from Figure 9(b), participants have the lowest app usage duration in MindShift (1.11 ±0.7 hours) compared to the Baseline (1.23±0.7 hours) and MindShift-Simple (1.20 ± 0.8 hours). Compared to Baseline, MindShift reduces by 9.8% usage duration while MindShift-Simple reduces by 2.4%. We conduct a Friedman test and find a significant difference among different techniques (p¡0.05). A post-hoc Wilcoxon test shows that MindShift has a trend of declining compared to Baseline, with marginal significance (p¡0.1).
(a) Total App Opening Frequency
(b) Total Phone Usage (Hours) per Day
Figure 10. App Usage Behavior
7.2.2. Habitual Usage Behavior
[…]
7.3. Subjective Report
We further analyze on user-reported SAS and self-efficacy scale results. When using MindShift and MindShift-Simple, participants experience a significant decrease in SAS Score and a significant increase in self-efficacy score, but they see no change when using Baseline. We summarize the results as follows.
7.3.1. Decrease of SAS Score
Both MindShift and MindShift-Simple decrease SAS Score significantly and MindShift achieves best. Figure 10(a) shows the results of the SAS scores during the intervention stages. MindShift exhibits the lowest SAS scores (35.2±10.2), followed by MindShift-Simple in the second position (37.6±10.1), with the Baseline intervention ranking the third (44.5±8.7) and the initial ranking the last (47.5±6.7). We conduct a Friedman test and the results show a significant difference(p¡0.001). Post-hoc Wilcoxon tests show that both MindShift and MindShift-Simple are significantly lower than the Baseline and initial scores (all ps¡0.01). This suggests that the two versions of MindShift have the potential to fundamentally alter individuals’ mobile phone usage behavior.
7.3.2. Increase of Self-efficacy
MindShift and MindShift-Simple increase Self-efficacy Score significantly. Figure 10(b) shows the results of the self-efficacy scores during the intervention stages. MindShift (26.7±6.1) and MindShift-Simple (26.6±6.8) exhibit higher scores compared to the Baseline intervention (24.1±5.6) and the initial (24.1±5.6). We conduct a Friedman test, and the results show a significant difference (p¡0.001). The post-hoc Wilcoxon test shows that both MindShift and MindShift-Simple are significantly higher than the Baseline and initial scores (all ps¡0.05). This indicates that persuasive techniques have the potential to significantly enhance individuals’ self-efficacy, which can result in overcoming excessive mobile phone usage. In contrast, conventional reminders may not achieve this goal.
(a) SAS Score
(b) Self-efficacy Score
Figure 11. Subjective Scales Report
7.3.3. Subjective Comments
During the exit interview, participants generally had a positive experience. One participant said, “I can actually feel that my recent dependency on the phone has decreased” (P17). Another participant felt using the app shifted them to self-improvement tasks instead of mindlessly using the smartphone, “Now, when I have nothing to do, I tend to do other things, like learning vocabulary, instead of aimlessly browsing my phone” (P10). Additionally, participants had positive comments on our design which includes the mental states factor during the intervention, “I feel that its suggestions align well with my emotional state at that time.” (P25). They also valued the “Understanding” and “Comforting” strategies, saying “It tells me that I’m not the only one experiencing these painful emotions, which helps me feel better” (P22).
Despite the majority positive comments, a small number of participants expressed their dissatisfaction with MindShift. Some participants found the persuasive message to be “a bit stiff and templated” (P4, P6), and they believed that “they would develop tolerance as they repeatedly use” (P20). Some participants also mentioned privacy concerns. For example, P15 mentioned that MindShift-Simple’s ability to capture time and location made her uncomfortable.
7.4. Summary of Results
Overall, two versions of MindShift show significantly higher acceptance rates compared to the Baseline. MindShift has the highest acceptance (35.2%) and thumb-up rates (6.8%), and it significantly outperforms MindShift-Simple in both acceptance rates of generated persuasion content (8.1%) and on a per-round basis (6-14.7%). Furthermore, there exist strategies that significantly outperform MindShift-Simple (3.6-10.8%) in every mental state, suggesting that the strategies we design are meaningful. Regarding app usage behavior, two versions of MindShift lead to a decrease in overall app opening frequency (12.1-14.4% ) and usage duration (2.4-9.8%) and are significantly effective in reducing habitual use. MindShift achieves the best results in reducing SAS scores (34.7%) and increasing self-efficacy scores (10.7%). Statistical analysis indicates that two versions of MindShift reduce SAS scores and increase self-efficacy scores significantly while Baseline (traditional pop-up reflections) does not. This suggests that MindShift has the potential to profoundly transform human behavior with enduring effects. Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and an inclination to continue using MindShift.
8. DISCUSSION
In this section, we discuss the MindShift’s novelty in contrast to previous intervention techniques and discuss the limitations and future work.
8.1. The Roles of Users’ Phone Use Purpose and Mental States in Smartphone Intervention
Most previous intervention techniques are designed to initiate interventions based on the amount of time and frequency of smartphone usage. However, quantifying smartphone use just by considering time oversimplifies and ignores the underlying causes. People use smartphones for work, study, and relaxation. MindShift initiates persuasion only when users agree the current usage belongs to habitual use. Our intervention enhances users’ self-awareness of their habitual phone use behavior.
Additionally, certain mental states are linked to habitual smartphone use as a form of self-distraction. Although smartphone use serves as a coping mechanism for emotion fluctuations, studies show that habitually using them for escapism fails to effectively mitigate emotions (Duvenage et al., 2020). Using smartphones for emotional regulation can lead to problematic smartphone use behavior, potentially leading to severe psychological issues such as depression (Coyne et al., 2021; Zsido et al., 2021). MindShift aims to intervene in habitual smartphone usage triggered by specific mental states. Our goal is to reduce users’ problematic smartphone use and help individuals transition from avoidance-oriented coping to approach-oriented coping (Compas et al., 2001). Future work can also explore automatic detection of users’ phone use intention and mental states, simplifying the intervention process and improving user experience. However, it remains an open question on how skipping the self-reflection may impact the effectiveness of such a persuasion technique.
8.2. Towards Adaptive Persuasion Intervention
MindShift generates dynamic and personalized persuasion content by combining information such as users’ physical contexts, mental states, and other behaviors. However, several participants still mentioned that the generated content sometimes could be “stiff and templated”. This may be attributed to the limited prompt templates. Although the content generated by the LLM varies, the main theme is guided by our prompts, which could limit the variation of persuasion content. To improve, we suggest to integrate user feedback into the system. Currently, MindShift supports a simple thumb-up and thumb-down feedback mechanism. Even with such simple information, we could establish a human-in-the-loop setup to fine-tune content, aligning better with user preferences. Moreover, future work can also consider collecting more comprehensive feedback from users. Users could customize the language style generated by an LLM, which can be coupled with adaptive algorithms such as reinforcement learning to achieve a more intelligent just-in-time adaptive intervention (JITAI) system that evolves with users  (Goldstein et al., 2017).
8.3. Ethical Concerns and Risk of Using LLM for Persuasion
Although MindShift performs well in reducing smartphone use frequency, there are important ethical concerns we want to highlight, especially about the application of LLMs for persuasion. Despite our intentionally devised prompts given to the LLM, it is not possible for us to guarantee that the generated persuasive messages are always safe. For example, if users access a blacklisted shopping app, LLM might include unwelcome commercial ideas in the message, causing user aversion and rejection. Meanwhile, because of LLM’s limited capability in understanding human psychology and society (Xu et al., 2023), it may evoke goals or recommend habits at the wrong time. For instance, if users are already stressed due to their life objectives (e.g., struggling with academic stress), additional reminders could exacerbate the stress or even cause harm. Recent research in the field of AI alignment aims to ensure that AI systems behave in ways that are safe, beneficial, and aligned with human values and objectives (Gabriel, 2020). Real-world deployment is needed to integrate these techniques into the intervention.
Meanwhile, privacy is another critical concern since our system detects users’ location (as general categories) and captures their smartphone usage data. MindShift employs a commercial API, transmitting users’ data to a third party. Although we intentionally designed the prompt to avoid including any identifiable information, there is still the risk of revealing information about their behavior and mental states. One solution for future study is to leverage open-sourced LLMs (such as LLaMA2 (Touvron et al., 2023) or PaLM2 (Anil et al., 2023)) so that user’s data can be appropriately handled and encrypted by ourselves instead of a third party.
8.4. Limitations
There are a few limitations in our study. First, our experimental user group is limited to college students, which limits the applicability of the results. Future studies can expand the sample size and involve more diverse user groups. Second, our field experiment is short. A five-week deployment cannot reveal the longitudinal effect of such an intervention technique. Moreover, if our time and monetary budget allow, our experiment design can be improved by making the Baseline another two-week intervention session for a more fair comparison. Third, our current method for detecting habitual use and mental states relies on self-reporting, which increases users’ burden. We only consider initial habitual use and do not consider the shifts in user purposes during app usage. As we mentioned in Section 8.1, future work can explore automatic intent and mental state detection, and the data collected in this study can serve as a starting point for machine learning training models. Finally, our study utilizes GPT-3.5 as a large language model, and its performance is still unstable. Future work can explore more lighted weighted and robust LLMs for local deployment, which can address the concerns mentioned in Section 8.3 to some extent.
9. CONCLUSION
In this paper, we propose MindShift, a mental-based persuasion intervention technique powered by LLM that aims to reduce problematic smartphone use. We conduct a Wizard-of-Oz study and an interview study to explore the mental states behind problematic smartphone use: stress, boredom, inertia, and design four persuasion strategies: understanding, comforting, evoking, and scaffolding habits. MindShift (1) collects users’ purpose, usage behavior, physical context, mental states, goals&habits, (2) uses the persuasion strategy we design, (3) leverages LLMs to generate dynamic, personalized persuasion messages. Through a five-week within-subjects user experiment (N=25), we compare three intervention techniques (MindShift, MindShift-Simple, Baseline). Our results show that MindShift significantly improves the acceptance rate of generated persuasion content (8.1%) compared with MindShift-Simple. Two versions of MindShift decrease overall app opening frequency (12.1-14.4%) and usage duration (2.4-9.8%) and are particularly effective in reducing habitual use. Moreover, two versions of MindShift significantly reduced SAS scores (25.8-34.7%) and increased self-efficacy scores (10.4-10.7%). Finally, users’ subjective comments also confirm a perceived reduction in smartphone dependency and a willingness to continue to use MindShift. Our work provides valuable insights into the mental states behind problematic smartphone use and the effectiveness of LLMs-powered persuasion for smartphone intervention.